OpenStudy (anonymous):

What is the fourth term in the expansion of (x + 2)7?

OpenStudy (anonymous):

What is the fourth term in the expansion of (x + 2)7? A) 280x4 B) 168x4 C) 320x2 D) 120x3

OpenStudy (anonymous):

@Not_Unknown Do you know the combination?

OpenStudy (anonymous):

Combination? The row? I know the row involves 1,7,21,35,35,21,7,1

OpenStudy (anonymous):

Yes, 7 C 3 = 35

OpenStudy (anonymous):

So the fourth term : = 7 C 3 * 2³ X⁴ = ....

OpenStudy (anonymous):

\[(x+2)^7=x^7+14 x^6+84 x^5+280 x^4+560 x^3+672 x^2+448 x+128 \]\[280 x^4 \]